# Fidelity Sweet Spot in Transmon Qubit Rings under Strong Connectivity Noise

### v1

@@ III.1 Sequence of SWAP operations
In quantum information processing, the SWAP gate plays a crucial role in enhancing connectivity beyond immediate neighbors, facilitating the transport of quantum information throughout the system. The duration of the interaction for one SWAP gate is given by $\tau=\frac{\pi}{4J}$, where $J$ represents the qubit-qubit coupling strength and can be identified as $J_{ij}$ in the control Hamiltonian (Eq. 1).
@@
First, we explore a relatively simple quantum circuit composed of a series of SWAP gates. As shown in Fig. 3, we select the initial state $|\Psi\rangle_{1}=\left|\uparrow\downarrow\downarrow\cdots\right\rangle$, which is an unentangled product state. The sequence of operations is as follows: starting from the state $\left|\Psi\right\rangle_{1}$, we swap the spin-up qubit all around the ring and then back again. We then numerically determine the fidelity of this sequence of operations, defined as the probability that the system, when measured, is found in the state corresponding to an error-free ($H_{\mathrm{noise}}=0$) execution of the intended sequence.
We calculate the infidelity of the circuit by varying the pulse time $\tau=\frac{\pi}{4J}$ for a given average noise level $\lambda_{0}$. For each value of $J/\lambda_{0}$, we generate 1,500 noise configurations sampled from the Gaussian distributions, compute the fidelity for each configuration, and then average over all samples to obtain the mean infidelity. The resulting dependence of the average infidelity on $J/\lambda_{0}$ is depicted in Fig. 4. It is natural to expect a higher fidelity with increased pulse strength (i.e., larger $J$) under a fixed noise level ($\lambda_{0}$), suggesting a reduced noise effect or, equivalently, achieving a higher $J/\lambda_{0}$ ratio should result in higher fidelity under constant pulse strength.
Figure 3: Quantum circuit implementing SWAP operations on a product state $|\Psi\rangle_{1}$. The left side represents the initial state, while the right side shows a sequence of SWAP gates transferring the qubit states along the chain.
In the context of ideal experimental conditions, where noise is minimal or pulse strength is significant, this monotonous relationship is also demonstrated in Ref. Foulk et al. (2022); Throckmorton and Das Sarma (2020). However, in real experiments, noise levels cannot always be minimized, and there is a limit to how high the pulse strength can be increased without introducing errors. The formula $\tau=\frac{\pi}{4J}$ demonstrates that higher pulse strengths result in shorter pulse times, which increases experimental complexity, indicating that the realistically achievable pulse strengths may not always optimize fidelity. Therefore, scenarios with lower pulse strengths, equivalently longer pulse times, or higher noise levels are more representative of actual experimental conditions.
In Fig. 4, the parameter region most indicative of real-world experiments is where $J/\lambda_{0}$ is less than $10^{2}$. An interesting result is observed within this range of higher noise or lower pulse strength, which is that the infidelity for all cases fluctuates with varying $J/\lambda_{0}$, exhibiting multiple minima. In the figure, we draw two horizontal lines at $\overline{F}=84.0\%$, 90.0% and 99.0%, the first two of which are comparable to typical experimental values Kandel et al. (2019); Sigillito et al. (2019). Notably, for the device with $L=4$, $\overline{F}=90\%$, the typical experimental fidelity Kandel et al. (2019), can be achieved near $J/\lambda_{0}=10$. This implies that the fidelity achievable under a 1 % noise level can also be attained at a 10 % noise level. Similarly, these sweet spots are observed for other system sizes, although they fall short of achieving high fidelity expected for quantum information processing.
Figure 4: Average infidelity of SWAP operations on the product state $|\Psi\rangle_{1}$ for different system sizes $L$. Horizontal dashed lines mark fidelity benchmarks of $\overline{F}=84\%$, $90\%$, and $99.9\%$. As $J/\lambda_{0}$ increases, the noise level decreases and fidelity improves monotonically. For $L=4$, the optimal performance, comparable to current experimental limits, is achieved near $J/\lambda_{0}\approx 10$.
III.2 The fidelity of general operations
Next, we extend our consideration to general quantum circuits. We characterize these operations using their distance to identity, defined as Nielsen (2002)
@@ III.2 The fidelity of general operations
where $d=2^{L}$ represents the dimension of the Hilbert space for the given systems. For each value of $D$, we randomly generate a unitary transformation $R$ with that distance to identity. To compute the fidelity for a given $R$, we average over 1,500 noise configurations sampled from the Gaussian distributions, following the same procedure as in Section III.1. This allows us to investigate whether sweet spots exist under different general quantum circuits and whether higher fidelity can be achieved at a lower experimental cost. The procedure to generate $R$ is detailed in Appendix B. A schematic figure depicting the evolution of $|\Psi\rangle_{1}$ under a randomly generated quantum circuit is shown in Fig. 5.
@@
Figure 5: Schematic of the product state $|\Psi\rangle_{1}$ evolved under a general random quantum circuit $R$. The left side shows the initial state $|\Psi\rangle_{1}$, while the right side represents the randomly generated operation acting on it.
@@
As illustrated in Fig. 6, sweet spots are present across different randomly generated quantum circuits. Each curve in the figure corresponds to a single $R$ with a specific $D$ value, and the fidelity is the average over 1,500 noise configurations. We observe a clear positive correlation between the distance $D$ and the fidelity: circuits closer to the identity (smaller $D$) yield higher fidelity. This is an expected result. However, more importantly, the fidelity sweet spots for different quantum circuits all occur at approximately the same $J/\lambda_{0}$ value. This indicates that once a sweet spot is measured for one circuit in a system, it can be used to optimize other circuits, enabling higher fidelity without additional experimental overhead.
@@
Figure 6: The fidelity of state $|\Psi\rangle_{1}$ for randomly generated quantum circuits with different distances $D$ to identity. Each curve corresponds to a single unitary operation $R$ with a specific $D$, and the fidelity is averaged over 1,500 noise configurations. Sweet spots are observed at similar $J/\lambda_{0}$ values across different circuits.
III.3 The impact of different states on fidelity
We next examine how different initial states affect the operation fidelity, focusing on the entangled states $|\Psi\rangle_{\mathrm{S}}$ and $|\Psi\rangle_{\mathrm{GHZ}}$. Since the main distinction among various quantum circuits lies in the magnitude rather than the position of the fidelity sweet spot, we restrict our analysis, without loss of generality, to a sequence of SWAP gates. For the state $|\Psi\rangle_{\mathrm{S}}$, the procedure is analogous to the product-state case: the spin on the right of the entangled pair is sequentially swapped around the ring and then brought back, effectively transferring entanglement from the first qubit to others along the chain. For the $|\Psi\rangle_{\mathrm{GHZ}}$ state, all qubits are initially entangled, and the operation corresponds to consecutive swaps between adjacent qubits. In a noisy system, such repeated exchanges are expected to cause stronger decoherence and entanglement loss, as each SWAP acts on a fully entangled configuration.
In Fig. 7, we evaluate the fidelities of the states $|\Psi\rangle_{\mathrm{S}}$ and $|\Psi\rangle_{\mathrm{GHZ}}$ under a sequence of SWAP operations in a system of size $L=6$. The initial states are assumed to be ideally prepared, free from any noise perturbations. This controlled setting enables a direct comparison of how noise influences the fidelity of these two types of entangled states during their evolution.
@@ III.3 The impact of different states on fidelity
Figure 7: Quantum circuits for preparing (a) $|\Psi\rangle_{\mathrm{S}}$ and (b) $|\Psi\rangle_{\mathrm{GHZ}}$ in a six-qubit ($L=6$) system. The shaded regions indicate the state-preparation stages, assumed to be ideally realized without noise, while the subsequent SWAP operations on the right are subject to connectivity-induced noise.
As shown in Fig. 8, all three initial states exhibit fidelity sweet spots under the sequence of SWAP operations. However, several additional observations emerge. First, the sweet spot positions differ among the states, indicating that even under fixed device parameters and identical circuits, the optimal operating point depends on the initial quantum state. More interestingly, contrary to the naïve expectation that repeated swaps of entangled states in a noisy system would reduce fidelity, we observe the opposite trend. The $|\Psi\rangle_{\mathrm{GHZ}}$ state achieves the highest overall fidelity, with several sweet spots approaching the ideal value of 99.9The fidelity of $|\Psi\rangle_{\mathrm{S}}$, involving a two-qubit entangled pair, is also superior to that of the unentangled product state $|\Psi\rangle_{1}$. This suggests that, within our system, a higher degree of initial entanglement can enhance robustness against connectivity noise. Nevertheless, the effect likely also reflects a symmetry correspondence between the initial state and the SWAP circuit, as the global symmetry of $|\Psi\rangle_{\mathrm{GHZ}}$ appears to align particularly well with that of the gate sequence. The detailed role of state-circuit symmetry in determining fidelity remains an open question Cabrera and Baylis (2007); Sudha et al. (2011). In our case, the superior performance of the highly entangled $|\Psi\rangle_{\mathrm{GHZ}}$ state highlights a practically relevant result for experiments using transmon qubit rings.
@@
Figure 8: Fidelity of different initial states under SWAP operations for $L=6$. Contrary to expectation, the fully entangled state $|\Psi\rangle_{\mathrm{GHZ}}$ attains the highest fidelity, reaching the ideal $\overline{F}=99.9\%$ at several sweet spots. The partially entangled state $|\Psi\rangle_{\mathrm{S}}$ also outperforms the unentangled product state $|\Psi\rangle_{1}$.
III.4 The origin of the fidelity sweet spot
@@
Our numerical results demonstrate that the quantum state fidelity in this system exhibits a characteristic sweet spot. This phenomenon has practical importance for real quantum circuits: under substantial noise or extended operation times, there exists an optimal evolution duration—or equivalently, a circuit depth—that maximizes fidelity. From an experimental perspective, the position of this optimum is more relevant than the exact fidelity value, since once the system parameters are fixed, the location of the fideility sweet spot remains essentially independent of the specific circuit.
@@
To uncover the physical origin of this effect, we analyze an analytically solvable two-qubit toy model with controlled approximations (see Appendix C). In this framework, the fidelity sweet spot of the $|\uparrow\downarrow\rangle$ state arises from the intrinsic oscillatory nature of quantum coherence: because $|\uparrow\downarrow\rangle$ is not an eigenstate of the noise Hamiltonian, its coherence exhibits temporal oscillations that periodically restore fidelity. By contrast, the GHZ state shows no such behavior in this simplified model, as the omission of superconducting-qubit frequency terms effectively makes it an eigenstate of the Hamiltonian, thereby suppressing coherence oscillations. When realistic frequency terms are included, we expect GHZ-type states to display analogous fidelity revivals and sweet-spot behavior.
@@
The universality of this mechanism comes from a general principle: any initial state that deviates from the Hamiltonian eigenbasis will exhibit oscillatory coherence dynamics under quasi-static noise. This feature is expected for systems dominated by slow noise fluctuations whose characteristic timescale exceeds that of the intrinsic quantum evolution Wendin (2017). In contrast, for other noise types—such as high-frequency broadband (Markovian), low-frequency narrowband (non-Markovian), or more complex forms including periodic, burst-like, or strongly correlated noise—the fidelity behavior becomes considerably richer. In these regimes, standard analyses based on simple coherence oscillations are inadequate, and approaches such as Lindblad master equations, hierarchical equations of motion (HEOM), Floquet theory, quantum-trajectory methods, or many-body simulations are required Gorini et al. (1976); Becker and Eckardt (2025); Zhang et al. (2024); Zhou et al. (2023); Carmichael (1993); Dalibard et al. (1992). Whether fidelity sweet spots persist under such generalized noise conditions remains an open question for future investigation.
IV A machine learning approach to predict the sweet-spot location
@@ IV.1 Machine learning framework
The numerical analyses above determine the fidelity sweet spot only for systems with fixed parameters. In experiments, however, both the CPW coupling strength and the noise distribution can vary and must often be measured individually. Because these factors directly influence the sweet-spot position—and thus the optimal circuit duration—performing exhaustive simulations or measurements for every configuration would be computationally and experimentally inefficient. To address this, we employ a supervised machine-learning strategy based on neural networks to predict the fidelity sweet spot of a given quantum device using only a limited set of simulation data. Once trained, this model can estimate the optimal circuit duration for new parameter regimes defined by different CPW coupling strengths and noise characteristics.
The overall workflow is as follows. We perform supervised learning by partitioning the available dataset into training and validation subsets, using the labeled data to fit the model and then testing its predictive performance on unseen examples. Each data point is characterized by an input vector comprising the system size $L$, the CPW noise strength $\lambda_{K}$, and the noise-distribution width $\sigma_{K}$, while the corresponding output vector contains the sweet-spot location and its associated fidelity. As the cost function—or empirical risk in decision theory—we adopt the mean squared error (MSE) loss Murphy (2012, 2022); Heaton (2018),
@@
where $n$ denotes the total number of data points, $y_{i}$ the observed (target) value for the $i$-th sample, and $\hat{y}_{i}$ the corresponding model prediction. This loss function is well suited for regression tasks, as it minimizes the mean squared deviation between predicted and true values. Although our model is nonlinear, we further evaluate its performance using the coefficient of determination,
$R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}},$ (6)
to assess the reliability of the training results. Here, $\bar{y}$ denotes the mean of the observed values and quantifies how well the model reproduces the data relative to their average. In practice, the MSE and $R^{2}$ metrics serve complementary roles: the MSE measures the absolute magnitude of prediction errors, whereas $R^{2}$ indicates the improvement over a baseline model that always predicts the mean. In our case, each $y_{i}$ corresponds to a two-dimensional vector $(J^{*}/\lambda_{0},1-\overline{F}^{*})$, representing the sweet-spot position and its associated infidelity. Together, these metrics provide a consistent framework for evaluating and optimizing the model’s performance. The overall machine-learning workflow is illustrated in Fig. 9.
@@
Figure 9: Illustration of the supervised-learning workflow using a multilayer-perceptron (MLP) neural network. The input parameters are the system size $L$, CPW noise strength $\lambda_{K}$, and noise-distribution width $\sigma_{K}$, while the outputs are the sweet-spot position $(J^{*}/\lambda_{0})$ and the corresponding infidelity $(1-\overline{F}^{*})$. The dataset is divided into training (80%) and validation (20%) subsets. For clarity, the diagram shows a simplified network with one hidden layer of six neurons; the actual model employs two hidden layers with ten neurons each, ReLU activations, and a linear output layer for two-dimensional regression.
@@
In supervised learning, the output of our model is a two-dimensional vector representing the sweet-spot location and the corresponding fidelity. To map these quantities, we employ a multi-layer perceptron (MLP) architecture, also known as a fully connected neural network Rumelhart et al. (1986); Hinton and Salakhutdinov (2006). The network comprises two hidden layers with ten neurons each, using the Rectified Linear Unit (ReLU) activation function Sudjianto et al. (2020). ReLU provides nonlinear expressivity at low computational cost, enabling efficient learning of complex relationships. The output layer uses a linear activation function, appropriate for regression tasks with continuous, real-valued outputs. In this case, the network produces a two-dimensional output vector, as illustrated in Fig. 9. The model is trained using the Adam optimizer Kingma and Ba (2017), an adaptive learning-rate algorithm that combines the advantages of Adagrad and RMSProp, and is well suited for general optimization tasks Duchi et al. (2011); Ruder (2016); Wilson et al. (2017).
IV.2 Data extraction and neural-network training

### v2

@@ III.1 Optimal gate duration of SWAP operations
In quantum information processing, SWAP gates play a crucial role in enhancing connectivity by enabling quantum state transport between non-adjacent qubits. The operation time for a single SWAP gate is given by $\tau=\pi/(4J)$, where $J$ corresponds to the qubit-qubit coupling strength $J_{ij}$ in the control Hamiltonian (Eq. 1).
Using the characteristic coupling strength $J\approx 10$ MHz from Table 1, the gate duration scales as $\tau\approx 80\,\text{ns}\times(100/J\lambda_{0}^{-1})$. Thus, our scanned range $J/\lambda_{0}\in[1,1000]$ corresponds to $\tau$ from $\sim 8\,\mu\text{s}$ down to $\sim 8\,\text{ns}$. The intermediate coupling regime ($J/\lambda_{0}\sim 10$100100), where optimal points appear, corresponds to $\tau\sim 80$$800\,\text{ns}$, which is longer than typical high-fidelity gate durations ($\sim 10$$50\,\text{ns}$) but still within practical coherence windows.
@@
We first investigate a quantum circuit composed of sequential SWAP operations, as illustrated in Fig. 3 for a representative system size of $L=6$. Starting from the unentangled product state $|\Psi\rangle_{1}=|\uparrow\downarrow\downarrow\cdots\rangle$, we perform a series of SWAP gates that transport the spin up qubit around the ring and back to its original position. The fidelity of this operation sequence is computed as the probability of obtaining the expected final state under ideal (noise free) evolution.
We systematically vary the ratio $J/\lambda_{0}$ from 1 to $10^{3}$, covering the weak to strong coupling regimes. Figure 4 shows the average infidelity as a function of $J/\lambda_{0}$ for different system sizes. For each $J/\lambda_{0}$ value, we sample 1,500 noise configurations from Gaussian distributions and compute the mean fidelity. In ideal conditions with minimal noise or strong pulses, fidelity typically improves monotonically with increasing $J/\lambda_{0}$, as demonstrated in previous studies Foulk et al. (2022); Throckmorton and Das Sarma (2020). However, practical experimental constraints—where noise cannot be arbitrarily suppressed and pulse strengths face technical limits—make the intermediate coupling regime ($J/\lambda_{0}<10^{2}$) particularly relevant.
Our results reveal a non-monotonic behavior in this experimentally significant regime. Rather than a steady improvement, the infidelity exhibits distinct minima—optimal operation points—where fidelity is maximized despite substantial noise levels. For the $L=4$ device, a fidelity of $\overline{F}=90\%$ (comparable to current experimental benchmarks Kandel et al. (2019); Sigillito et al. (2019)) is achieved near $J/\lambda_{0}=10$. This demonstrates that performance typically expected under 1% noise levels can be attained even with 10% noise through optimal gate duration selection. Similar optimal points appear for larger system sizes, though the achievable fidelity decreases with increasing qubit number.
Figure 3: Quantum circuit implementing SWAP operations on product state $|\Psi\rangle_{1}$ for system size of $L=6$ qubits as a representative example. The sequence transfers qubit states around the ring, with gate duration determined by the coupling strength $J$.
[…]
III.2 Optimal gate duration of general operations
We now extend our analysis to general quantum operations beyond simple SWAP gates. To characterize the complexity of these operations, we employ the distance to identity metric Nielsen (2002):
@@ III.2 Optimal gate duration of general operations
where $d=2^{L}$ is the Hilbert space dimension. For each distance value $D$, we randomly generate a unitary transformation $R$ with that specific distance to identity. The generation procedure is detailed in Appendix B.
Figure 5 illustrates the evolution of the product state $|\Psi\rangle_{1}$ under a general random quantum circuit for $L=6$. For each operation $R$, we compute the fidelity by averaging over 1,500 noise configurations following the same methodology as in Section III.1.
@@
Figure 5: Evolution of product state $|\Psi\rangle_{1}$ under a general random quantum circuit $R$ for $L=6$ qubits. The left panel shows the initial state, while the right panel depicts the action of the randomly generated operation.
@@
As shown in Fig. 6 (for $L=4$), optimal operation points emerge consistently across different randomly generated quantum circuits. Each curve corresponds to a specific unitary $R$ with a particular distance $D$, with fidelity values representing averages over noise configurations. We observe a clear correlation between the distance $D$ and the achievable fidelity: operations closer to the identity (smaller $D$) yield higher fidelity, as expected. More significantly, however, the optimal operation points for different quantum circuits all occur at approximately the same $J/\lambda_{0}$ value. This consistency indicates that once an optimal operating point is identified for one circuit in a given system, it can be leveraged to optimize other circuits without additional experimental overhead, enabling enhanced fidelity across multiple operations.
@@
Figure 6: Fidelity of state $|\Psi\rangle_{1}$ for randomly generated quantum circuits with varying distances $D$ to identity for $L=4$ qubits. Each curve represents a single unitary operation $R$ with specific $D$, with fidelity averaged over 1,500 noise configurations. Optimal operation points occur at similar $J/\lambda_{0}$ values across different circuits.
III.3 Impact of initial states on optimal points
We next investigate how different initial states affect operation fidelity, focusing on the entangled states $|\Psi\rangle_{\mathrm{S}}$ and $|\Psi\rangle_{\mathrm{GHZ}}$. Since the primary distinction among quantum circuits lies in fidelity magnitude rather than optimal point position, we restrict our analysis to SWAP gate sequences without loss of generality.
Figure 7 illustrates the quantum circuits for preparing and evolving these entangled states in a six qubit system ($L=6$). For the singlet state $|\Psi\rangle_{\mathrm{S}}$, we sequentially swap the right qubit of the entangled pair around the ring and back, effectively distributing entanglement along the ring. For the GHZ state $|\Psi\rangle_{\mathrm{GHZ}}$, where all qubits are initially entangled, we perform consecutive swaps between adjacent qubits. The initial state preparation (shaded regions) is assumed ideal, enabling direct comparison of noise effects during the SWAP operations.
(a)(b)
@@ III.3 Impact of initial states on optimal points
Figure 7: Quantum circuits for (a) $|\Psi\rangle_{\mathrm{S}}$ and (b) $|\Psi\rangle_{\mathrm{GHZ}}$ in a six-qubit system. Shaded regions indicate ideal state preparation, while subsequent SWAP operations are subject to connectivity noise.
As shown in Fig. 8, all three initial states exhibit clear optimal operation points under SWAP sequences. However, several important observations emerge. First, the positions of these optimal points vary among states, indicating that even with fixed device parameters and identical circuits, the optimal operating condition depends on the initial quantum state. More remarkably, contrary to the expectation that entangled states would suffer greater decoherence under repeated noisy operations, we observe the opposite trend: the fully entangled GHZ state achieves the highest overall fidelity, with several optimal points approaching the quantum error correction threshold of 99.9%. The partially entangled singlet state $|\Psi\rangle_{\mathrm{S}}$ also outperforms the unentangled product state $|\Psi\rangle_{1}$.
This suggests that higher initial entanglement can enhance robustness against connectivity noise in our system. The superior performance of the GHZ state may reflect favorable symmetry alignment between the initial state and the SWAP circuit structure, though the precise role of state-circuit symmetry requires further investigation Cabrera and Baylis (2007); Sudha et al. (2011). These results have practical implications for experimental implementations using transmon qubit rings, indicating that carefully chosen initial states can significantly improve operational fidelity.
@@
Figure 8: Fidelity of different initial states under SWAP operations for $L=6$. The fully entangled GHZ state achieves the highest fidelity, reaching 99.9% at several optimal points, while the singlet state outperforms the product state.
III.4 Origin of the fidelity optimal operation point
@@
Our numerical results demonstrate that the quantum state fidelity in transmon qubit rings exhibits characteristic optimal operation points. This phenomenon has practical significance for real quantum circuits: under substantial noise or extended operation times, there exists an optimal evolution duration—equivalently, an optimal gate duration—that maximizes fidelity. From an experimental perspective, the position of this optimum is more critical than the exact fidelity value, as it remains largely independent of the specific circuit once system parameters are fixed.
@@
To elucidate the physical origin of this effect, we analyze an analytically solvable two qubit model with controlled approximations (see Appendix C). In this framework, the fidelity optimal point for the $|\uparrow\downarrow\rangle$ state arises from the intrinsic oscillatory nature of quantum coherence. Since $|\uparrow\downarrow\rangle$ is not an eigenstate of the noise Hamiltonian, its coherence exhibits temporal oscillations that periodically restore fidelity. In contrast, the GHZ state shows no such behavior in this simplified model due to the omission of superconducting qubit frequency terms, which effectively renders it an eigenstate of the Hamiltonian, thereby suppressing coherence oscillations. When realistic frequency terms are included, we expect GHZ type states to display analogous fidelity revivals and optimal point behavior.
@@
The universality of this mechanism stems from a general principle: any initial state that deviates from the Hamiltonian eigenbasis will exhibit oscillatory coherence dynamics under quasistatic noise. This characteristic is typical for systems dominated by slow noise fluctuations whose timescale exceeds that of the intrinsic quantum evolution Wendin (2017). For other noise types—such as high frequency broadband (Markovian), low frequency narrowband (non-Markovian), or more complex forms including periodic, burst ike, or strongly correlated noise—the fidelity behavior becomes considerably richer. In these regimes, standard analyses based on simple coherence oscillations are inadequate, and approaches such as Lindblad master equations, hierarchical equations of motion (HEOM), Floquet theory, quantum trajectory methods, or many body simulations are required Gorini et al. (1976); Becker and Eckardt (2025); Zhang et al. (2024); Zhou et al. (2023); Carmichael (1993); Dalibard et al. (1992). Whether fidelity optimal points persist under such generalized noise conditions remains an open question for future investigation.
IV A machine learning approach to predict the optimal operation point
@@ IV.1 Machine learning framework
The numerical analyses presented thus far determine optimal operation points only for systems with fixed parameters. In experimental settings, however, both CPW coupling strengths and noise distributions can vary significantly and must often be characterized individually. Since these factors directly influence the optimal operation point—and consequently the ideal circuit duration—performing exhaustive simulations or measurements for every configuration would be computationally and experimentally prohibitive.
To address this challenge, we employ a supervised machine learning approach based on neural networks to predict the optimal operation point for a given quantum device using only limited simulation data. Once trained, this model can estimate the optimal circuit duration for new parameter regimes defined by different CPW coupling strengths and noise characteristics.
The overall workflow, illustrated in Fig. 9, proceeds as follows. We perform supervised learning by partitioning the available dataset into training and validation subsets, using labeled data to fit the model and then testing its predictive performance on unseen examples. Each data point is characterized by an input vector comprising the system size $L$, CPW noise strength $\lambda^{K}$, and noise distribution width $\sigma_{K}$, while the corresponding output contains the optimal operation point position $(J^{*}/\lambda_{0})$ and its associated infidelity $(1-\overline{F}^{*})$.
As the cost function, we adopt the mean squared error (MSE) loss Murphy (2012, 2022); Heaton (2018):
@@
where $n$ denotes the total number of data points, $y_{i}$ the observed value, and $\hat{y}_{i}$ the model prediction. We further evaluate performance using the coefficient of determination:
$R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}$ (6)
where $\bar{y}$ denotes the mean of observed values. In our case, each $y_{i}$ corresponds to a two-dimensional vector $(J^{*}/\lambda_{0},1-\overline{F}^{*})$. These complementary metrics provide a consistent framework for evaluating model performance.
@@
Figure 9: Supervised learning workflow using a multilayer perceptron neural network. Input parameters are system size $L$, CPW noise strength $\lambda^{K}$, and noise distribution width $\sigma_{K}$; outputs are the optimal operation point position and corresponding infidelity. The dataset is divided into training (80%) and validation (20%) subsets. For clarity, the diagram shows a simplified network with one hidden layer of six neurons; the actual model employs two hidden layers with ten neurons each, ReLU activations, and a linear output layer for two dimensional regression.
@@
We employ a multilayer perceptron (MLP) architecture with two hidden layers of ten neurons each, using Rectified Linear Unit (ReLU) activations for nonlinear expressivity Sudjianto et al. (2020). The output layer uses linear activation appropriate for regression tasks. The model is trained using the Adam optimizer Kingma and Ba (2017), which combines advantages of Adagrad and RMSProp for efficient optimization Duchi et al. (2011); Ruder (2016); Wilson et al. (2017).
IV.2 Data extraction and neural network training